# Partial Resolution of the Erdős-Straus, Sierpiński, and Generalized Erdős-Straus Conjectures Using New Analytical Formulas

### v1

@@ 7 Second Formula: Condition of Existence of a Perfect Square
Let $y=t(4x-n)-q$ and $z=t(4x-n)+q$. We need to show that, with this formula, the conjecture holds for those $q$ such that
@@
$=\frac{1}{x}+\frac{4x-n}{nx}=\frac{n-n+4x}{nx}=\frac{4}{n}.$
Thus, the second formula allows us to affirm that for any $n\in\mathbb{N}$ with $n\geq 2$, if there exists $(x_{0},t_{0})\in(\mathbb{N}^{*})^{2}$ such that
@@
then there exist $y=t_{0}(4x_{0}-n)-q$ and $z=t_{0}(4x_{0}-n)+q$ with $(y,z)\in\mathbb{N}^{2}$, such that
@@ 8.1 Transforming the Equation
For the right-hand side to be strictly positive, it is necessary that $05x-n>$ 0, i.e., $x>\frac{n}{5}$, leading to $x\geq\left\lfloor\frac{n}{5}\right\rfloor+1$.
@@
$y+z=(5x-n)(2t),\quad yz=nx(2t),\quad\forall t\in\mathbb{N}^{*}.$ (6)
@@ 8.2 Special Case: $n$ Divisible by 3
8.3 General Case: Finding $x$ Values for which $\Delta>0$
Expanding $\Delta$:
$\Delta=100t^{2}x^{2}-40t^{2}xn+4t^{2}n^{2}-8ntx.$ (10)
Solving for $x$ such that $\Delta>0$, we obtain:
$x\in\left]0,\frac{n(5t+1-\sqrt{10t+1})}{25t}\right[\cup\left]\frac{n(5t+1+\sqrt{10t+1})}{25t},+\infty\right[.$ (11)
@@ 8.3 General Case: Finding $x$ Values for which $\Delta>0$
we can set $t=k=k(5x-n)^{2}-2nx$, ensuring that $k\in\mathbb{N}^{*}$ if $15x-n-$ 1 and $15x-n+$ 1 divide $2nx$.
Thus, we find:
@@ 8.4 Conclusion
To conclude, for any $n\geq 2$, if there exists $(x_{0},t_{0})\in\mathbb{N}^{*2}$ such that:
@@
then $y=t_{0}(5x_{0}-n)-q$ and $z=t_{0}(5x_{0}-n)+q$ satisfy the equation, proving the conjecture under this condition.
@@ 9 Generalization of the Erdős-Straus and Sierpiński Conjectures
We recall that the generalized of the Erdős-Straus and Sierpiński Conjectures states that for any nonzero natural number $a$ with $a\geq 4$, and for any natural number $n$, $n\geq 2$, there exist $x,y,z$ in $\mathbb{N}^{*}$ such that:
$\frac{a}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$
It should be noted that since the inception of this conjecture, there has not been significant research on it, and thus no concrete leads have emerged. In this paper, I will attempt to partially prove it as I did with the two conjectures above, and I will reduce the difficulty to verifying two conditions to completely settle all three conjectures. Here, I will use the same approach as in the previous two conjectures.
@@ 9.1 Trivial Case
As before, let us start with the trivial case: If $(a-2)$ divides $n$, then we can simply take $x=\frac{n}{a-2}$ and $y=z=n$, which proves the conjecture in this particular case.
@@ 9.2 Equivalence of the conjecture
We will show that the generalized Erdős-Straus and Sierpiński conjecture is true if and only if $y$ and $z$ are solutions of the following quadratic equation in $V$:
@@
That is, we will establish that the equation
@@
is equivalent to finding $y,z$ as solutions of
$V^{2}-(ax-n)(2t)V+2nxt=0.$
9.3 Preliminary Theorem
Before proving this equivalence, let us state the following theorem:
Theorem (Vieta’s Formulas – Equivalence Between Coefficients and Roots of a Quadratic Polynomial)
Let $X_{1}$ and $X_{2}$ be two real (or complex) numbers, and consider the quadratic equation:
@@ 9.3 Preliminary Theorem
where $S$ and $P$ are two real(or complex) parameters.
Then, the following two statements are equivalent:
1. $X_{1}$ and $X_{2}$ are solutions of the quadratic equation above.
2. $X_{1}$ and $X_{2}$ satisfy the relations:
$X_{1}+X_{2}=S,\quad X_{1}\cdot X_{2}=P.$
@@ 2. $X_{1}$ and $X_{2}$ satisfy the relations:
9.4 Proof of the Equivalence
Let $a\in\mathbb{N}$ with $a\geq 4$, $n\in\mathbb{N}$ with $n\geq 2$, and $x,y,z\in\mathbb{N}^{*}$. We have:
@@ 9.4 Proof of the Equivalence
For $\frac{y+z}{yz}>0$, it is necessary that $ax-n>0$, i.e., $x\geq\left\lceil\frac{n}{a}\right\rceil+1$.
@@
for any $t\in\mathbb{N}^{*}$.
By substituting $y+z$ with $(ax-n)(2t)$ and $yz$ with $nx(2t)$ in $yz(ax-n)=nx(y+z)$, the equation remains valid, which shows that this equation implies the system $(S)$.
@@ 9.5 Reciprocal Proof
for any $t\in\mathbb{N}^{*}$, then it follows that $yz(ax-n)=nx(y+z)$.
Indeed, since $y+z=(ax-n)(2t)$, we obtain:
@@
noting that $ax-n\neq 0$ since we imposed $x\geq\left\lceil\frac{n}{a}\right\rceil+1$ to ensure $ax-n>0$.
@@
$nx(y+z)=yz(ax-n).$
This proves that system $(S)$ implies $nx(y+z)=yz(ax-n)$.
From the previous paragraph, we have justified the equivalence between $nx(y+z)=yz(ax-n)$ and the system $(S)$.
Finally, by applying Vieta’s theorem stated earlier, we conclude that the system $(S)$ is equivalent to stating that $y$ and $z$ are solutions of the following quadratic equation in $V$:
@@ 10 Summary of the generalized Erdős-Straus and Sierpiński conjecture
always has solutions in $\mathbb{N}^{*}$ for any $a\in\mathbb{N}$ with $a\geq 4$ and any $n\in\mathbb{N}$ with $n\geq 2$. This reduces the difficulty of proving the conjecture to finding integer solutions (in $\mathbb{N}^{*}$) for a second-degree equation, whose discriminant is given by:
@@
Thus, the proof of the generalized Erdős-Straus and Sierpiński conjecture can be reduced to verifying the following two conditions or refuting one of them to disprove the conjecture:
  1. 1.
    $t^{2}(ax-n)^{2}-2ntx$ is always a perfect square for all $(a,n)$.
  2. 2.
    $02t(ax-n)\pm\sqrt{\Delta}>$ 0. because condition (1) is sufficient, to obtain integer solutions, since $2t(ax-n)$ and $\sqrt{\Delta}$ will always be even, if (1) is verified, and the factor 2 in the formula will be simplified during the calculation of :
    $V=\frac{2t(ax-n)\pm\sqrt{\Delta}}{2}$ (23)
    is valid.
@@
$x\in\left]0,\frac{n(at+1-\sqrt{2at+1})}{a^{2}t}\right[\cup\left]\frac{n(at+1+\sqrt{2at+1})}{a^{2}t},+\infty\right[\quad\text{with }x\in\mathbb{N}.$ (27)
similary by doing the same manipulations for the x for which $\Delta>0$
@@
Verifying,
@@
Similarly, the second formula allows us to state that for any $a\in\mathbb{N}$ with $a\geq 4$ and for any $n\in\mathbb{N}$ with $n\geq 2$, if there exist $(x_{0},t_{0})\in\mathbb{N}^{*2}$ such that:
@@ 11 Comparison with Mordell’s Approach
Mordell’s method provided a partial resolution of the Erdős-Straus conjecture by demonstrating that for all integers $n$ such that $n\not\equiv 1,11^{2},13^{2},17^{2},19^{2},23^{2}\pmod{840}$, the conjecture holds. However, for the exceptional cases left unresolved by Mordell, our second formula proves to be a powerful tool for generating solutions.
Through numerical testing, we observed that by naively programming the second formula, we can capture a significant portion of the exceptional cases left by Mordell. Specifically:
  • Among the values of $n$ in the range $[2,2000]$, we capture 86% of the cases not covered by Mordell.
11.1 Testing Second Formula for Mordell’s Exceptional Cases between [2, 2000]
illustration (Here, $L$ calculates the ratio $\frac{4}{n}$ and $R$ calculates the sum $\left(\frac{1}{x}+\frac{1}{y}+\frac{1}{z}\right)$, and the code outputs true if $R=L$ and false otherwise). Mordell’s exceptional cases in the given range : $121,169,289,361,529,841,961,1009,1129,1201,1369,1681,1801,1849,1969$ Total exceptional cases: 15
For $n=121$ (Mordell case), $x=33$, $y=726$, $z=726$, $t=66$: $L=0.0331$, $R=0.0331$, Conjecture holds: True
For $n=169$ (Mordell case), $x=44$, $y=1690$, $z=2860$, $t=325$: $L=0.0237$, $R=0.0237$, Conjecture holds: True
For $n=289$ (Mordell case), $x=85$, $y=578$, $z=2890$, $t=34$: $L=0.0138$, $R=0.0138$, Conjecture holds: True
For $n=361$ (Mordell case), $x=92$, $y=8664$, $z=10488$, $t=1368$: $L=0.0111$, $R=0.0111$, Conjecture holds: True
For $n=529$ (Mordell case), $x=138$, $y=6348$, $z=6348$, $t=276$: $L=0.0076$, $R=0.0076$, Conjecture holds: True
For $n=841$ (Mordell case), $x=220$, $y=6380$, $z=18502$, $t=319$: $L=0.0048$, $R=0.0048$, Conjecture holds: True
For $n=961$ (Mordell case), $x=248$, $y=15376$, $z=15376$, $t=496$: $L=0.0042$, $R=0.0042$, Conjecture holds: True
For $n=1009$ (Mordell case), $x=5676$, $y=264$, $z=3818056$, $t=88$: $L=0.0040$, $R=0.0040$, Conjecture holds: True
For $n=1129$ (Mordell case), $x=3952$, $y=304$, $z=2230904$, $t=76$: $L=0.0035$, $R=0.0035$, Conjecture holds: True
For $n=1369$ (Mordell case), $x=348$, $y=38332$, $z=45066$, $t=1813$: $L=0.0029$, $R=0.0029$, Conjecture holds: True
For $n=1681$ (Mordell case), $x=423$, $y=121032$, $z=138744$, $t=11808$: $L=0.0024$, $R=0.0024$, Conjecture holds: True
For $n=1849$ (Mordell case), $x=473$, $y=40678$, $z=40678$, $t=946$: $L=0.0022$, $R=0.0022$, Conjecture holds: True
For $n=1969$ (Mordell case), $x=495$, $y=177210$, $z=177210$, $t=16110$: $L=0.0020$, $R=0.0020$, Conjecture holds: True
Percentage of Mordell’s exceptional cases verified: 86.67%
  • In the range $[2000,5000]$, we capture 60% of the remaining cases.
11.2 Testing Second Formula for Mordell’s Exceptional Cases: between [2000, 5000]
Mordell’s exceptional cases in the given range:
$\begin{array}[]{c}2041,2209,2521,2641,2689,2809,2881,3049,3361,3481,\\ 3529,3649,3721,3889,4201,4321,4369,4489,4561,4729\\ \end{array}$
Total exceptional cases: 20
Testing Second Formula for Mordell’s Exceptional Cases:
$\begin{array}[]{|c|c|c|c|c|c|c|c|}\hline\cr n&x&y&z&t&L&R&Conjectureholds\\ \hline\cr 2041&520&48984&61230&1413&0.0020&0.0020&True\\ 2209&564&53016&53016&1128&0.0018&0.0018&True\\ 2641&665&184870&184870&9730&0.0015&0.0015&True\\ 2809&710&112360&150520&4240&0.0014&0.0014&True\\ 2881&731&97954&97954&2278&0.0014&0.0014&True\\ 3481&885&104430&104430&1770&0.0011&0.0011&True\\ 3649&984&21894&29192&89&0.0011&0.0011&True\\ 3721&976&29768&59536&244&0.0011&0.0011&True\\ 4201&1056&369688&403296&16804&0.0010&0.0010&True\\ 4321&1085&483952&503440&25984&0.0009&0.0009&True\\ 4369&1104&200974&209712&4369&0.0009&0.0009&True\\ 4489&1139&152626&152626&2278&0.0009&0.0009&True\\ \hline\cr\end{array}$
Percentage of Mordell’s exceptional cases verified: 60.00%
In the range $[5000,10000]$, we capture 61% of the remaining cases.
11.3 Testing Second Formula for Mordell’s Exceptional Cases : between [5000, 10000]
Mordell’s exceptional cases in the given range:
$\begin{array}[]{c}5041,5161,5209,5329,5401,5569,5881,6001,6049,6169,\\ 6241,6409,6721,6841,6889,7009,7081,7249,7561,7681,\\ 7729,7849,7921,8089,8401,8521,8569,8689,8761,8929,\\ 9241,9361,9409,9529,9601,9769\end{array}$
Total exceptional cases : 36
Testing Second Formula for Mordell’s Exceptional Cases:
$\begin{array}[]{|c|c|c|c|c|c|c|c|}\hline\cr n&x&y&z&t&L&R&Conjectureholds\\ \hline\cr 5041&1278&181476&181476&2556&0.0008&0.0008&True\\ 5161&1314&115924&185796&1588&0.0008&0.0008&True\\ 5329&1460&21316&53290&73&0.0008&0.0008&True\\ 5401&1353&1328646&1328646&120786&0.0007&0.0007&True\\ 6001&1765&12002&60010&34&0.0007&0.0007&True\\ 6049&1518&798468&798468&34716&0.0007&0.0007&True\\ 6169&1544&2680384&2763712&388864&0.0006&0.0006&True\\ 6241&1580&249640&249640&3160&0.0006&0.0006&True\\ 6409&1612&512720&548080&13600&0.0006&0.0006&True\\ 6721&1683&2056626&2056626&186966&0.0006&0.0006&True\\ 6889&1743&289338&289338&3486&0.0006&0.0006&True\\ 7009&1763&574738&574738&13366&0.0006&0.0006&True\\ 7081&1848&74690&96360&275&0.0006&0.0006&True\\ 7249&1815&2392170&2392170&217470&0.0006&0.0006&True\\ 7729&1947&510114&510114&8646&0.0005&0.0005&True\\ 7849&1974&659316&659316&14028&0.0005&0.0005&True\\ 7921&2136&47526&63368&89&0.0005&0.0005&True\\ 8401&2108&1142536&1142536&36856&0.0005&0.0005&True\\ 8569&2145&3341910&3341910&303810&0.0005&0.0005&True\\ 9361&2343&3987786&3987786&362526&0.0004&0.0004&True\\ 9409&2716&18818&263452&97&0.0004&0.0004&True\\ 9529&2932&19058&38116&13&0.0004&0.0004&True\\ \hline\cr\\ \end{array}$
Percentage of Mordell’s exceptional cases verified: 61.11%
These results show that the second formula is a machine for generating solutions, including for the cases left open by Mordell(These results show that this formula naively captures 69% of the cases of $n$ in the range $[2,10000]$ that satisfy the conjecture but are not captured by Mordell’s formula). Moreover, even with a relatively simple computational implementation, a large number of solutions have been found, suggesting that a complete proof of the conjecture could be within reach.
Finally, it is worth noting that these statistics can be further refined by increasing the precision of the search for optimal parameters $x$ and $t$. This work thus opens up a new and promising perspective in the search for a complete proof of the Erdős-Straus conjecture.
To allow further verification and exploration beyond the values presented in this article, we include the implementation of our computational method in the appendix. This will enable researchers to test additional values of $n$ without overloading the main content of this paper.

### v2

@@ 7 Second Formula: Condition of the Existence of a Perfect Square
Let y=t(4xn)qy=t(4x-n)-q and z=t(4xn)+qz=t(4x-n)+q. We need to show that, with this formula, the conjecture holds for those $q$ such that
@@
$\frac{1}{x}+\frac{4x-n}{nx}=\frac{n-n+4x}{nx}=\frac{4}{n}.$
Thus, the second formula allows us to state the following theorem.
Theorem 2.
For any $n\in\mathbb{N}$ with $n\geq 2$, if there exists $(x_{0},t_{0})\in(\mathbb{N}^{*})^{2}$ such that
@@
then there exist
$y=t_{0}(4x_{0}-n)-q\quad\text{and}\quad z=t_{0}(4x_{0}-n)+q,$
with $(y,z)\in\mathbb{N}^{2}$, such that
@@ 8.1 Transforming the Equation
For the right-hand side to be strictly positive, it is necessary that 5xn>05x-n>0, i.e., $x>\frac{n}{5}$, leading to $x\geq\left\lfloor\frac{n}{5}\right\rfloor+1$.
@@
$y+z=(5x-n)(2t),\quad yz=nx(2t),\quad\ t\in\mathbb{Q}_{+}^{*}.$ (6)
@@ 8.2 Special Case: $n$ Divisible by 3
8.3 First Formula for the Sierpiński Conjecture
Solving for $t$ such that $\Delta>0$, we obtain:
$t>\frac{2nx}{(5x-n)^{2}}.$
@@ 8.3 First Formula for the Sierpiński Conjecture
we can set $t=k=k(5x-n)^{2}-2nx$, Thus, we find:
@@ 8.4 Conclusion
We may therefore conclude this section with the following theorem.
Theorem 3.
For any integer $n\geq 2$, if there exists $(x_{0},t_{0})\in\mathbb{N}^{*2}$ such that
@@
then the quantities
$y=t_{0}(5x_{0}-n)-q\quad\text{and}\quad z=t_{0}(5x_{0}-n)+q$
satisfy
$\frac{5}{n}=\frac{1}{x_{0}}+\frac{1}{y}+\frac{1}{z},$
with $(y,z)\in\mathbb{N}^{*}$.
@@ 9 Generalization of the Erdős-Straus and Sierpiński Conjectures
We recall that the generalized of the Erdős-Straus and Sierpiński Conjectures states that for every integer $a\geq 5$, there exists an integer $N_{0}\in\mathbb{N}$ with $N_{0}\geq 2$ such that for all $n\geq N_{0}$, there exist $x,y,z\in\mathbb{N}^{*}$ satisfying
$\frac{a}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$ (18)
It should be noted that since the inception of this conjecture, there has not been significant research on it, and thus no concrete leads have emerged. In this paper, we will attempt to partially prove it as I did with the two conjectures above, and we will reduce the difficulty to verifying two conditions to completely settle all three conjectures. Here, we will use the same approach as in the previous two conjectures.
@@ 9.1 Trivial Case
As before, let us start with the trivial case: If (a2)(a-2) divides $n$, then we can simply take $x=\frac{n}{a-2}$ and $y=z=n$, which proves the conjecture in this particular case.
@@ 9.2 Equivalence of the Conjecture
Theorem 4.
Let $a\geq 4$ be a fixed integer. There exists an integer $N_{0}\in\mathbb{N}$ with $N_{0}\geq 2$ such that, for all $n\geq N_{0}$, the generalized Erdős–Straus and Sierpiński conjecture holds if and only if there exist $x\in\mathbb{N}^{*}$ and $t\in\mathbb{Q}_{+}^{*}$ such that $y$ and $z$ are solutions of the quadratic equation in $V$:
@@
That is,we will establish that for all $n\geq N_{0}$, the identity
@@
is satisfied if and only if $y$ and $z$ are roots of the quadratic equation above.
9.3 Preliminary Result
Proposition 1 (Vieta’s Formulas).
Let $X_{1}$ and $X_{2}$ be two real (or complex) numbers, and consider the quadratic equation
@@ 9.3 Preliminary Result
where $S$ and $P$ are real (or complex) parameters.
Then the following statements are equivalent:
  1. 1.
    $X_{1}$ and $X_{2}$ are solutions of the quadratic equation.
  2. 2.
    $X_{1}$ and $X_{2}$ satisfy the relations
    $X_{1}+X_{2}=S\quad\text{and}\quad X_{1}X_{2}=P.$
@@
9.4 Proof of the Theorem (Equivalence)
Let $a\in\mathbb{N}$ with $a\geq 4$, $n\in\mathbb{N}$ with $n\geq N_{0}$, and Suppose that there exist integers $x,y,z\in\mathbb{N}^{*}$ such that:
@@ 9.4 Proof of the Theorem (Equivalence)
For $\frac{y+z}{yz}>0$, it is necessary that axn>0ax-n>0, i.e., $x\geq\left\lceil\frac{n}{a}\right\rceil+1$.
@@
for $t\in\mathbb{Q}_{+}^{*}$.
By substituting $y+z$ with (axn)(2t)(ax-n)(2t) and $yz$ with $nx(2t)$ in yz(axn)=nx(y+z)yz(ax-n)=nx(y+z), the equation remains valid, which shows that this equation implies the system $(S)$.
@@ 9.5 Reciprocal Proof
for $t\in\mathbb{Q}_{+}^{*}$, then it follows that yz(axn)=nx(y+z)yz(ax-n)=nx(y+z).
Indeed, since y+z=(axn)(2t)y+z=(ax-n)(2t), we obtain:
@@
noting that $ax-n\neq 0$ since we imposed $x\geq\left\lceil\frac{n}{a}\right\rceil+1$ to ensure axn>0ax-n>0.
@@
nx(y+z)=yz(axn).nx(y+z)=yz(ax-n).
This proves that system $(S)$ implies nx(y+z)=yz(axn)nx(y+z)=yz(ax-n).
From the previous paragraph, we have justified the equivalence between nx(y+z)=yz(axn)nx(y+z)=yz(ax-n) and the system $(S)$.
Finally, by applying Vieta’s formulas given in the above proposition, stated earlier, we conclude that the system $(S)$ is equivalent to stating that $y$ and $z$ are solutions of the following quadratic equation in $V$:
@@ 10 Summary of the generalized Erdős-Straus and Sierpiński conjecture
always has solutions in $\mathbb{N}^{*}$ for any $a\in\mathbb{N}$ with $a\geq 4$ and any $n\in\mathbb{N}$ with $n\geq N0$. This reduces the difficulty of proving the conjecture to finding integer solutions (in $\mathbb{N}^{*}$) for a second-degree equation, whose discriminant is given by:
@@
$t>\frac{2nx}{(ax-n)^{2}}.$ (24)
similary by doing the same manipulations for the t for which $\Delta>0$
@@
The quantities $y$ and $z$ are integers if and only if $2nx$ is divisible by both axn1ax-n-1 and axn+1ax-n+1. in this case $x$,$y$,$z$ satisfy:
@@
Similarly, the second formula allows us to state that for any $a\in\mathbb{N}$ with $a\geq 4$ and for any $n\in\mathbb{N}$ with $n\geq N0$, if there exist $(x_{0},t_{0})\in\mathbb{N}^{*2}$ such that:
@@ 11 Comparison with Mordell’s Approach
Mordell’s method provided a partial resolution of the Erdős–Straus conjecture by showing that for all integers $n$ such that
$n\not\equiv 1,11^{2},13^{2},17^{2},19^{2},23^{2}\pmod{840}$,
@@
the conjecture holds. However, for the exceptional cases left unresolved by Mordell, our second formula proves to be a powerful tool for generating solutions.
Through numerical testing, we observed that a direct implementation of the second formula, applied to the values of $n$ not covered by Mordell’s result in the range $2\leq n\leq 10\,000$, successfully produces solutions for all such $n$. In this interval, this corresponds to a numerical success rate of $100\%$.
Remark: This observation is purely experimental and does not constitute a proof. However, it strongly suggests that our approach captures the full underlying structure of the original Erdős–Straus equation. This is very likely due to the equivalence established earlier between the original Diophantine equation and our associated quadratic parametric equation, from which integer solutions are obtained directly via the analysis of its discriminant.
Moreover, since the Erdős–Straus conjecture has already been verified by other authors for all $n\leq 10^{17}$, it follows from the proven equivalence that our quadratic equation must necessarily admit integer solutions at least up to this bound in the classical case $a=4$. From this perspective, it is therefore not surprising that our formula successfully produces solutions for the values of $n$ left unresolved by Mordell, in particular for all $n$ in the interval [2,10 000][2,10\,000].
For completeness and reproducibility, we provide in the appendix the numerical codes used to test the cases not covered by Mordell’s result, as well as to explore all values of $n$ in user-specified intervals. These codes apply more generally to all fixed integers $a\geq 4$, including the classical Erdős–Straus case ($a=4$), the Sierpiński case ($a=5$), and their natural generalizations ($a\geq 6$).
12 Testing Second Formula for Mordell’s Exceptional Cases between [2,5000]
Mordell’s exceptional cases in the given range are:
[…]
Total exceptional cases: $35$
$n$ $x$ $t$ $y$ $z$
121 31 1078 1694 4774
169 44 325 1690 2860
289 75 510 2550 8670
361 92 1368 8664 10488
529 133 15778 42826 51842
841 220 319 6380 18502
961 248 496 15376 15376
1009 260 1009 10090 52468
1129 2964 104 312 2230904
1201 9548 775 310 57335740
1369 348 1813 38332 45066
1681 423 11808 121032 138744
1801 35720 342 456 96497580
1849 468 3483 60372 99846
1969 493 215818 633302 661606
2041 516 4239 73476 121518
2209 564 1128 53016 53016
2521 18908 163 652 23833534
2641 665 9730 184870 184870
2689 9412 181 724 12654434
2809 710 4240 112360 150520
2881 724 18963 242004 326886
3049 75768 275 770 165011880
3361 73525 34 850 19769402
3481 885 1770 104430 104430
3529 27056 342 912 71610468
3649 915 55432 569244 650260
3721 943 2806 115046 171166
3889 40836 83 996 26468534
4201 1056 16804 369688 403296
4321 1085 25984 483952 503440
4369 1104 4369 200974 209712
4489 1139 2278 152626 152626
4561 15964 307 1228 36405902
4729 244728 66 1188 128590968
The above table shows that the second formula successfully produces valid integer solutions for all Mordell exceptional values of $n$ in the interval $[2,5000]$. Consequently, a numerical success rate of $100\%$ is achieved on this entire set of previously unresolved cases.
13 Testing Second Formula for Mordell’s Exceptional Cases between [5000,10000]
Mordell’s exceptional cases in the given range are:
[…]
Total exceptional cases: $36$.
$n$ $x$ $t$ $y$ $z$
5041 1264 56800 806560 897440
5161 1314 1588 115924 185796
5209 96888 275 1320 210287330
5329 1460 73 21316 53290
5401 1353 120786 1328646 1328646
5569 1112412 41 1394 364413084
5881 92628 83 1494 60527252
6001 1520 3177 180030 321936
6049 1518 34716 798468 798468
6169 1544 388864 2680384 2763712
6241 1580 3160 249640 249640
6409 1612 13600 512720 548080
6721 1683 186966 2056626 2056626
6841 32923 328 1804 81900452
6889 1743 3486 289338 289338
7009 1755 203424 2186808 2288520
7081 1815 1056 96360 281688
7249 1815 217470 2392170 2392170
7561 368600 190 1900 557396920
7681 47047 286 2002 103248002
7729 1935 247328 2658776 2782440
7849 1974 14028 659316 659316
7921 2136 89 47526 63368
8089 139536 342 2052 376235568
8401 2108 36856 1142536 1142536
8521 332320 268 2144 707924680
8569 2145 303810 3341910 3341910
8689 2184 17378 729876 903656
8761 2208 8761 403006 841056
8929 426921 88 2244 298978636
9241 181935 104 2340 149445452
9361 2343 362526 3987786 3987786
9409 2425 970 94090 470450
9529 2398 11728 639176 838552
9601 327639 62 2418 161316002
9769 2460 9769 586140 801058
The above table shows that the second formula successfully produces valid integer solutions for all Mordell exceptional values of $n$ in the interval $[5000,10000]$. Consequently, a numerical success rate of $100\%$ is achieved on this entire set of previously unresolved cases.
Taken together, the two preceding tables show that our second formula captures $100\%$ of the integers left unresolved by Mordell for all values of $n$ in the interval $[2,10000]$.
Any reader wishing to verify the correctness of these numerical simulations may proceed as follows: for a given value of $n$ and the corresponding triple $(x,y,z)$ provided in the tables, it suffices to check the identity
$4xyz=n(xy+xz+yz),$
which is strictly equivalent to the original Erdős–Straus equation
$\frac{4}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$
For large values of $n$, it is strongly preferable to verify solutions using the polynomial identity $4xyz=n(xy+xz+yz)$ rather than the fractional equation itself. Indeed, the latter involves floating-point arithmetic and may suffer from numerical round-off errors, potentially preventing exact equality from being detected by the machine. In contrast, the polynomial form relies solely on integer arithmetic and therefore provides an exact, unambiguous mathematical verification.